Link to report from morning meeting
1:13 - 2:27 stable beams Fill 2670, 1.5e11 average
bunch intensity, Peak lumi: 6.6e33 cm-2s-1, Integrated lumi ~ 27 pb-1,
Emittance in collision ~ 2.2 μm rad
2:27 Fill 2670 dumped by power converter problem of RQTX2.L8 (Free wheel
5:33 stable beams Fill 2671, 1.5e11 average bunch intensity, Peak lumi:
6.4e33 cm-2s-1, Emittance in collision ~ 2.3 μm rad
Cryo recovery Pt 8
22:00 cryo recovered
22:42 – 23:14 problem with RF module 2B2
Dry dump to verify LBDS
Fill 2668 dumped by losses at TCLA.C6R7.B1 at the very beginning of
collisions BP (bunch intensity a bit too high? average 1.53e11)
7B2 low level problem
Fill 2669 stable beams, average bunch intensity 1.5e11, peak lumi:
Beams dumped by operation due to cryo problem:
Refrigeration unit (QURC) down in point 8. Integrated lumi: 132 pb-1.
Access to US85 to reset the magnetic axle bearing of QURCB.
ALICE access (dipole and compensator had been ramped down before).
End of ALICE access.
rise in the triplet zone 7.4L1.X (no pump in the segment and triplet
- Cryo recovery ongoing.
- 1:06 Fill 2662 dumped
in adjust, high beam losses (bunches at the end of trains), average
bunch intensity 1.55e11.
- 3:44 – 15:04 stable
beams Fill 2663, 1.51e11 average bunch intensity, Peak lumi:
6.5e33 cm-2s-1, Integrated lumi 167 pb-1
– 15:00 tune scan B2
beam dumped for scheduled access: Atlas gas leak; LHCb; Pt 6 interlock
BPMS cabling check and attenuators; Vacuum Pt 4 to repair a pumping
group at 14R4; TOTEM LVDT bypass activated and B2 injection BIC channel
OFSU, LBDS and MKI problems
tripped due to a spike in U_res
- 00:41 Dumped by
interlocked BPM in IR6.
- 01:22 Problems with
TCL.5R1.B1 (recurrent issue). Disabled the resolver check on lost steps.
- 02:11 Injection.
- 03:53 Stable beams.
Peak lumi: 6.0e33.
- 05:14 RB.A45 QPS trip.
- 09:00 Access for IR8
triplet. Other accesses in shadow.
BPM interlock test
Fill 2660 stable beams (no more 'private' IP8 bunches), 1.48e11 average
bunch intensity, peak lumi 6.1e33 cm-2s-1, integrated lumi 24.5 pb-1
beam dumed by SIS BLM HV interlock (communication problem between BLM
CPU and CMW).
for Atlas gas leak
- 04:24 Operator beam
for physics. Delays due to problems with SPS bunch length.
Fill 2653 stable beams, 1.40e11 ppb, 5.6 1033 cm-2s-1peak
lumi, 11.6 pb-1 total lumi delivered.
Dumped by collimator interlock on TCDD (injection movable
mask/collimator) for beam 1.
Access for replacement of power supply for PXI rack.
Access finished. Recovering TCDD and TDI positions for beam 1. In
shadow: QPS board swap on RQTL11.L5B1 and RCD.A45B1.
Injection. Steering TI2 and TI8,
problems with BPM interlocks from LSS4 and LSS6. PS and SPS need to
readjust the beam.
TOTEM interlock for injection.
Injecting slightly higher bunch intensities: 1.47e11 average. Emittance from SPS: ~1.8 micron (on 12 b).
Stable beams. Peak lumi: 5.6e33.
- 00:42 Collimator
issues: RBAC access problem.
- 01:09 FBCT issue:
- 01:27 Interlock from
TCL.5R1.B1: missing steps.
- 02:49 Again FBCT
- 03:05 Injection
- 04:03 Interlock from
- 06:19 Injection.
- 08:09 Ramp.
- 08:48 stable beams
fill 2651. Peak Luminosity: 5.8 1033 cm-2s-1.
- Access for switching
temperature probe DFB L7.
- In shadow: Access for
fire detection equipment at point 1. ATLAS access.
- 12:26 Injection.
Interlock from TCL.5R1.L1: lost steps.
- 13:41 Stable beams.
5.9e33 lumi. 123MJ. Low losses in intensity.
- 20:59 End of fill. Trip
of IT.R1. Validation of inner triplet in R1 needed due to spurious
signals in PM. ATLAS intervention.
- 23:33 Triplet check
complete. Signals are normal and observed spike is due to a slight
current imbalance between the Q1 and the Q3 of the triplet. Proceed.